If the sides of a square are decreased by 3 cm, the area is decreased by 81 cm^2. What were the dimensions of the original square?

Thank you.

Answer by stanbon(75887)

Let the original square have side=x, and area=x^2
If the side is decreased by 3cm it becomes x-3, and area is (x-3)^2.
EQUATION:
old area - new area = 81 cm^2
x^2-(x-3)^2=81
x^2-[x^2-6x+9]=81
6x-9=81
6x=90
x=15cm (size of old side)
x-3=12cm (size of new side)
Cheers,
Stan H.